# Cartan–Thullen theorem and Levi problem in context of generalised convexity

### v2

@@ 5. Proofs
Let us now show that (v) is equivalent to (i). Note that any $\mathcal{F}$-polytope is $\mathcal{F}$-convex by Lemma 4.6. If $K\subset\Omega$ is compact, then it is contained in some $\mathcal{F}$-polytope, and so does its $\mathcal{F}$-convex hull. Thus, (i) follows. The converse implication follows thanks to an application of the equivalence of (i) and (ii). ∎
5.3. Proof of Theorem 1.23
Proof of Theorem 1.23.
We shall employ Theorem 1.11. Let $\mathcal{A}$ denote the space of real-parts of functions in $\mathcal{B}$. Proposition 4.9 shows that holomorphic convexity of $\Omega$, i.e., compactness of $\mathrm{clConv}_{\mathcal{A}}K$ for any compact $K\subset\Omega$ is equivalent to (i) of Theorem 1.11.
@@
Note that $\Omega$ is complete with respect to $\mathcal{A}$ is and only if it is complete with respect to $\mathcal{B}$. Similarly, it is an $\mathcal{A}$-space if and only if it is a $\mathcal{B}$-space. ∎

### v3

@@ 5 Proofs
Let us now show that (v) is equivalent to (i). Note that any $\mathcal{F}$-polytope is $\mathcal{F}$-convex by Lemma 4.8. If $K\subset\Omega$ is compact, then it is contained in some $\mathcal{F}$-polytope, and so does its $\mathcal{F}$-convex hull. Thus, (i) follows. The converse implication follows thanks to an application of the equivalence of (i) and (ii).
5.3 Proof of Theorem 1.23
Proof 5.15 (Proof of Theorem 1.23).
We shall employ Theorem 1.11 letting $\mathcal{A}$ denote the space of real-parts of functions in $\mathcal{B}$. Proposition 4.13 shows that holomorphic convexity of $\Omega$, i.e., compactness of $\mathrm{clConv}_{\mathcal{A}}K$ for any compact $K\subset\Omega$ is equivalent to (i) of Theorem 1.11.
@@
Note that $\Omega$ is complete with respect to $\mathcal{A}$ is and only if it is complete with respect to $\mathcal{B}$. Similarly, for any metric $d$, it is an $(\mathcal{A},d)$-space if and only if it is a $(\mathcal{B},d)$-space.
This shows that all conditions of Theorem 1.23 are equivalent.
6 The lemma of Bremermann and Lelong implies the resolution of the Levi problem
The lemma of Bremermann and of Lelong [4, Theorem 2] states that if $\Omega\subset\mathbb{C}^{n}$ is a domain of holomorphy, then any plurisubharmonic function on $\Omega$ is a Hartogs function. The set of Hartogs functions on $\Omega$, cf. [4, Section 2], is the smallest convex cone of functions on $\Omega$ that contains the logarithms of moduli of holomorphic functions on $\Omega$ and is closed under taking:
  1. (i)
    locally bounded suprema,
  2. (ii)
    monotonically decreasing limits.
Moreover, the function belongs to the set of Hartogs functions on $\Omega$ if and only if its restrictions to any open, precompact subset $U$ of $\Omega$ is a Hartogs function on $U$, and any upper semi-continuous regularisation of a Hartogs function is again a Hartogs function, that is if $u\colon\Omega\to\mathbb{R}$ is a Hartogs function then also so is
$z\mapsto\limsup_{z^{\prime}\to z}u(z^{\prime}).$
Remark 6.1.
We note that the proof of [4, Theorem 2] shows that any plurisubharmonic function $u\colon\Omega\to\mathbb{R}$ on a pseudoconvex domain $\Omega\subset\mathbb{C}^{n}$ can be written as
$u(z)=\limsup_{z^{\prime}\to z}\limsup_{m\to\infty}\frac{\log\lvert f_{m}(z^{\prime})\rvert}{m}$
for some sequence $(f_{m})_{n=1}^{\infty}$ of holomorphic functions on $\Omega$. We refer also to [2, Lemma 6.2].
Theorem 6.2.
The Bremermann–Lelong lemma is equivalent to the positive resolution of the Levi problem.
Proof 6.3.
If the Levi problem has positive resolution, then [4, Theorem 2] shows that the lemma holds true.
Suppose that the converse is valid. Let $\Omega\subset\mathbb{C}^{n}$ be pseudoconvex. We shall show that it is holomorphically convex. By Theorem 1.23 we see that $\Omega$ admits a continuous plurisubharmonic exhaustion function $p\colon\Omega\to\mathbb{R}$, which is a Hartogs function thanks to our assumption. More specifically, see Remark 6.1, it can be written as
$p(z)=\limsup_{z^{\prime}\to z}\limsup_{m\to\infty}\frac{\log\lvert f_{m}(z^{\prime})\rvert}{m}$
for some sequence $(f_{m})_{n=1}^{\infty}$ of holomorphic functions on $\Omega$.
Let $K\subset\Omega$ be compact. Since $\Omega$ is open, there exists $\epsilon>0$ such that the set
$K_{\epsilon}=\{\omega\in\Omega\mid\mathrm{dist}(\omega,K)\leq\epsilon\}$
is a compact subset of $\Omega$. Since $p$ is continuous, it is bounded by $C$ on $K_{\epsilon}$. Therefore for any $z\in K$
$\limsup_{m\to\infty}\frac{\log\lvert f_{m}(z)\rvert}{m}\leq C.$
Let $\mathcal{B}$ be the space of real-parts of holomorphic functions on $\Omega$ and let $\mathcal{S}$ be the convex cone generated by the logarithms of absolute values of holomorphic functions on $\Omega$. The set
$\mathrm{clConv}_{\mathcal{B}}K=\{\omega\in\Omega\mid a(\omega)\leq\sup a(K)\text{ for all }a\in\mathcal{B}\}$
is contained in the set
$L=\{\omega\in\Omega\mid b(\omega)\leq\sup b(K)\text{ for all }b\in\mathcal{S}\}.$
If $\omega\in L$, i.e., $b(\omega)\leq\sup b(K)$ for all $b\in\mathcal{S}$, then by the Hahn–Banach theorem for convex cones, [34, Theorem 2.1], and by the Riesz’ representation theorem of non-negative functionals there exists a probability measure $\mu_{\omega}$ on $K$ such that for all $b\in\mathcal{S}$
[…]
Therefore for all $m=1,2,\dotsc$ we have
[…]
Thus
[…]
Now, thanks to the boundedness of $p$ and thus also of $\limsup_{n\to\infty}\frac{1}{m}\log\lvert f_{m}\rvert$ on $K$ we can employ the Fatou lemma and see that
$p(\omega)=\limsup_{m\to\infty}\frac{1}{m}\log\lvert f_{m}\rvert(\omega)\leq\Bigg{(}\int_{K}\ \limsup_{m\to\infty}\frac{1}{m}\log\lvert f_{m}\rvert(\omega)\,d\mu_{\omega}\Bigg{)}\leq C.$
This completes the proof, as $p$ has compact sublevel sets and therefore $\mathrm{clConv}_{\mathcal{B}}K$ is compact for any compact $K\subset\Omega$. This is to say, $\Omega$ is holomorphically convex.